A man completes 5/8 of a job in 10 days. At this rate, how many more days will it take to finish the job?
Answer by TimothyLamb(4379)

rate = (5/8)/10 = 0.0625 jobs/day
days/job = 1/0.0625 = 16
---
Answer:
6 more days to finish
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php